Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (GMaccaferri at Fri May 29 13:32:57 METDST 1998) Success
Bad weather condition. Rain all the time

Torun (K. Borkowski at Fri May 29 13:45:40 METDST 1998) Success
Field System reported "unlocked" BBC2.

Westerbork (Tony Foley at Sat May 30 13:59:47 METDST 1998) Minor failures
Data looks OK, but we have been having decoder problems so I am looking at the eye pattern and searching on the oscilloscope for a synch pattern to check that there is data on the tape.

Onsala (A. Polatidis at Sun May 31 14:02:24 METDST 1998) Success
No known problems

Jodrell (Alastair Gunn at Mon Jun 1 15:40:43 METDST 1998) Minor failures
Diode cal drive in wrong position until 09:39 UT. Expect reduced signal-to-noise. No other known problems.

Effelsberg (dgraham at Tue Jun 2 18:31:27 METDST 1998) Minor failures
off source 0800-0912 du to op error

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:23:27 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name v123a2.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:14:16 METDST 1998) Success
Only 7 min lost for antenna in prelimit from 04:03 to 04:10